# Application of Analytical Hierarchical Process and its Variants on Remote Sensing Datasets

### v1

@@ Abstract.
Artificial Intelligence, Ganges, River Taxonomy, Remediation
copyright: none
1. Introduction
@@
The acute and chronic impacts of human activities on the river Ganga are physical, chemical, and biological (Simon and Joshi, 2022(24, ); Modi et al., 2023(12, )). For instance, the Ganga Basin is physically transformed with several barrages (i.e. dams) as well as an extensive irrigation system that diverts the Himalayan discharge away from river Ganga to support agriculture through the basin. Moreover, changes in water chemistry along the river are driven by a variety of factors, including industrial waste and synthetic fertilizer runoff, along the length of the river (Sharma et al., 2023(23, )). Additionally, Srinivas et al. (2020)(26, ) identified 13 threats and challenges impacting the river Ganga that included broad concerns ranging from biodiversity loss to open defecation near the river. To reduce anthropogenic induced threats and increase the overall health of the river basin, river rejuvenation and development must simultaneously address multiple social, economic, and environmental dimensions (Pal, 2023)(16, ).
@@
In order to prioritize areas of the Ganga basin for actions such as conservation, restoration, intervention, and rejuvenation, it is crucial to perform vulnerability assessments. There are many vulnerability assessment methods, and a multitude of factors, dimensions, and subdimensions may be included in an assessment (Plummer et al., 2012)(17, ). Critically, our assessment considers not only the immediate vulnerability of a given area but also the threat to the broader river system posed by potential sources of pollution in that area. This approach properly recognizes that the interaction of pollutants with local geophysical factors creates variation in environmental impacts downstream.
@@
The Analytical Hierarchical Process (AHP) is a multicriteria decision making process that combines pairwise comparisons and expert judgements to arrive at its conclusions. Developed by TL Saaty in 1980 (20, ), AHP has been used in numerous applications, ranging from resource allocation and site prioritization to risk assessment and conflict resolution. Climate and geophysical models have proven well-suited for AHP, due to the complexity of decision making under uncertainty with respect to factors impacting their features of interest. For example, Jhariya et al. (2021)(8, ) used AHP in combination with Geospatial Information Systems (GIS), remote sensing, and vertical electrical soundings (VES) to identify potential zones of groundwater using weighted data on geology, geomorphology, rainfall, lineament, LULC (land use and land cover), drainage density, slope, soil type, and soil texture. Using AHP, they were able to estimate the likelihood of groundwater in each zone, with five levels of potential (low, medium, medium-high, high, and very high) with an accuracy of 80%.
1.1. AHP factors of waterway vulnerability and their Saaty rankings
@@
In this study, AHP was used as a characterization scheme rather than a decision making process; hence, factors and their relative rank order of importance were structured to highlight areas that may be vulnerable to pollution with an emphasis on urban pollution. Land use was identified using LULC, a widely recognized direct characterization of land types and includes development such as urbanization and agricultural uses. Urbanization may be directly related to major sources of contamination in water bodies, while agriculture is a major source of nonpoint source pollution including fertilizers and pesticides. Moreover, densely populated areas have a direct impact on surface water quality. For instance, population growth rate and water quality parameters, such as biochemical oxygen demand (BOD) and dissolved oxygen, were highly correlated in a Kelani river watershed in Sri Lanka (Liyanage & Yamada, 2017)(9, ). Given our emphasis on urban pollution and their direct impacts on vulnerability to pollution, population density (PD) was ranked as the most important factor (1 on the Saaty-scale), and LULC was ranked second with a Saaty-scale value of 2 in relative importance.
@@
Rainfall, slope, and drainage density indirectly impact vulnerability in that they are not sources of pollution, yet they directly influence erodibility and pollution mobility. For example, the runoff from steeper slopes was more likely to carry pollutants into streams than runoff from land use in flatter slopes (Yu et al., 2016)(30, ), and effects on water quality originating from land use (agriculture, industrial, and residential) adjacent to water bodies were dependent on rainfall variability (Nobre et al., 2000)(15, ). Drainage density (DD) is an indicator of surface runoff processes and direct runoff and pollution transport increase with greater DD (Bedient et al., 1978)(3, ). Rainfall, DD, and slope were assigned Saaty-scale values of 4, 5, and 7 respectively because rainfall initiates pollutant movement events, slope is an important factor in pollutant transport, and DD influences final distribution of pollutants. Temperature is an environmental determinant of microbial activity and productivity including nuisance algal blooms. For instance, water surface temperature behavior in zones with algal bloom occurrences presented greater significant values, up to 3°C, than those with clearer water (Ferral et al., 2021)(6, ). Moreover, seasonal increases of anaerobic bacteria in aquatic surface sediment during summer compared to cool months indicated a depletion of oxygen in the overlying water (Smoot and Findlay, 2001)(25, ). Given the absence of high resolution site monitoring throughout the Ganga basin, land surface temperature (LST) was used as a surrogate temperature estimate. LST was assigned the lowest importance among all input factors (9 on the Saaty-scale) because it was expected to have a minor overall relative impact on vulnerability.
1.2. Potential Shortcomings of AHP
@@
While AHP offers numerous advantages in terms of decision making in remote sensing, Munier and Hontoria (2021)(13, ) identified 30 potential shortcomings of the method. Given the importance of accurately understanding vulnerability, we reviewed all the shortcomings of AHP and assessed whether they were applicable for remote sensing data and our particular use case (Supplemental Table S1). For instance, the first critique “The Pair-Wise Method and Its Application in AHP” suggested that some problems may not be suited for AHP if their decision space is amorphous with many interrelated factors that have minimal differences in their relative importance. This critique was not applicable for our use-case, since our problem fits well to the AHP architecture. Similarly, the critique around rank reversals, which arise when uncertainty in the decision space is larger than the consistency of rank, was applicable on our use-case since a change in parameter weights might lead to an inaccurate assessment of the vulnerability of the river. To test for this and design a better method, we applied AHP while taking into account the effect of unknown variables (1-N AHP) and Fuzzy AHP as diagnostic analyses. To summarize, all the applicable shortcomings were addressed with the following approaches: Analytic Network Process (ANP) (Saaty, 2006)(22, ), Fuzzy AHP (Saaty, 1982)(21, ), AHP with non-linear parametric influence (Nested AHP), 1-N AHP, and Fuzzy AHP while taking into account the effect of unknown factors (Fuzzy 1-N AHP) (Supplemental Table S1).
@@
In this paper, we apply AHP to our remote sensing datasets to assess the vulnerability of the river Ganga to pollution loads along a 1,330 km stretch of the river. We cover the problem formulation, results and inferences of AHP. We also include a deep-dive on each shortcoming of AHP in the context of Climate Modelling, addressing the list set forth by Munier and Hontoria (2021)(13, ). We share whether each shortcoming was applicable to our problem statement, while suggesting and attempting alternative solutions to counteract the shortcomings. Through this dialogical process, we have arrived at two novel decision making approaches inspired by AHP, namely Nested AHP and 1-N AHP. The methodological contributions of this research are, at one level, these new variants of AHP we propose; and at another level, the demonstration of how incremental evaluation of the AHP process and its shortcomings relative to a given problem can reveal application-specific permutations of the original. Finally, we report the pollution vulnerability evaluation of the river Ganga produced by these processes.
2. Methods and Data
2.1. Study area and data source
@@
The area of analysis (26,609.4 km2) is a 20 km zone extending 10 km on both sides of the main channel of river Ganga from Haridwar, Uttarakhand, where it makes its debouche from the Siwalik Hills into the Indo-Gangetic plains to its confluence with river Ghaghara at Sitab Diara, Uttar Pradesh, India.
@@
The change in elevation along this 1,330 km stretch of Ganga is from 314 m to 76 m above mean sea level. Data from Landsat-8 (Level 2, Collection 2, Tier 1; pixel resolution, 30x30m) and other sources (Supplemental Table S2) were used to assess vulnerability in the area of analysis include slope, drainage density (DD), land surface temperature (LST), rainfall, land use/land cover (LULC), and population density (Figure 2). The drainage density and center line for the river Ganga were both derived from a digital elevation model (DEM) of the watershed using typical GIS hydrography processes, as previously described by Rahaman et al. (2015)(19, ). Streams were extracted from NASADEM data as a vector format Shapefile using the R programming language, Whitebox, Terra, and sf packages. Lower order streams were removed to obtain the main Ganga river centerline with ArcGIS Pro. The Ganga center line down-river of the Uttar Pradesh border was removed, and the 20 km zone was established as a buffer around this shortened centerline within the river Ganga.
Figure 1. Inset Map of India, Individual States, Ganga Basin, and Our Scope
@@
Figure 1. This map shows the Ganga Basin, in gray, and the buffer region, in blue, which is the scope of our paper. The 20 km buffer zone extends 10 km on both sides of the main channel of river Ganga.
Figure 2. AHP: Individual Parameter Vulnerability Map
Figure 2. Visual representation of the data used in the calculation of AHP. Fig. 1a, 1c, 1d, 1e, and 1f are broken up into extreme low to extreme high using Natural Breaks. Fig. 1b is separated out by its unique values.
2.2. Analytical Hierarchy Process (AHP)
@@
To categorize the vulnerability of the area immediately around the river Ganga and to identify potential locations to implement our remediation strategies, we applied the AHP on Remote Sensing and other geospatial data as previously described by Jhariya et al. (2021)(8, ). Environmental, topographical, and anthropogenic factors were subjectively ranked from equal importance to relatively greatest hierarchical importance across the full breadth of the 9-point Saaty scale based on their likelihood of impacting vulnerability (Section 1.1). A pairwise matrix constructed from these ranked factors and their reciprocal values were used to estimate a mean normalized weight for each factor (equation (1)). These normalized weights were applied to each of five factor classes that were identified by natural breaks in the data using the Jenks method (Table 1). LULC categories were numerically scored as vulnerability criteria based on their linkage to pollution. Lastly, overall vulnerability per pixel was determined from the sum of the factor vulnerability scores. Additional details describing factor rank orders, pairwise comparison matrices and their weights, as well as Jenks natural breaks are provided in Supplemental Information.
@@
Table 1. Factor classes used in Analytical Hierarchy Process decision matrix ranked by vulnerabilitya
@@
5: Extremely High ¿4036 Built-up ¿1410 ¿4.0 ¿35 ¿37
@@
1: Extremely Low ¡865 Water/Wetland ¡868 ¡0.6 ¡4 ¡25
aFactor classes were based on natural breaks (Jenks and Caspall, 1971)(7, ).
@@ 2.2.1. Ranking method
(1) $A=\begin{bmatrix}1&2&4&5&7&9\\ 0.5&1&3&4&5&7\\ 0.25&0.33&1&2&4&5\\ 0.2&0.25&0.5&1&3&5\\ 0.14&0.2&0.25&0.33&1&4\\ 0.11&0.14&0.2&0.2&0.25&1\end{bmatrix}$
@@
(2) $\tilde{a}_{ij}=\frac{a_{ij}}{\sum_{i=1}^{n}a_{ij}}$
The priority vector $w_{AHP}$ (the eigenvector) was then derived by averaging the rows of the normalized matrix:
(3) $w_{i}=\frac{1}{n}\sum_{j=1}^{n}\tilde{a}_{ij}$
@@
We then checked the consistency of the pairwise comparison matrix by calculating the Consistency Index (CI) and the Consistency Ratio (CR) (Saaty, 1980)(20, ). The CI was calculated using the formula:
(4) $CI=\frac{\lambda_{max}-n}{n-1}$
@@
(5) $CR=\frac{CI}{RI}$
In our case, with a consistency ratio (CR) of 0.06, the matrix was considered consistent (as CR ¡ 0.1). This consistency indicated that our pairwise comparison matrix was reliable for determining the weights. To create the vulnerability map of the Ganga buffer region, if
@@
    (6) $W=\begin{bmatrix}w_{1}&w_{2}&w_{3}&w_{4}&w_{5}&w_{6}\end{bmatrix}$
@@
    (7) $X(p)=\begin{bmatrix}x_{1}(p)&x_{2}(p)&x_{3}(p)&x_{4}(p)&x_{5}(p)&x_{6}(p)\end{bmatrix}$
We calculated the predicted vulnerability for each pixel by multiplying each factor’s parameter value by its corresponding AHP weight i.e.:
(8) $V(p)=W\cdot X(p)^{T}$
@@
2.3. Diagnostic Analyses
2.3.1. Nested AHP

### v2

# SSupplemental References

@@ Abstract
keywords:
[…]
1 Introduction
@@
The acute and chronic impacts of human activities on the river Ganga are physical, chemical, and biological (Simon and Joshi, 2022; Modi et al., 2023). For instance, the Ganga Basin is physically transformed with several barrages (i.e. dams) as well as an extensive irrigation system that diverts the Himalayan discharge away from river Ganga to support agriculture through the basin. Moreover, changes in water chemistry along the river are driven by a variety of factors, including industrial waste and synthetic fertilizer runoff, along the length of the river (Sharma et al., 2023). Additionally, Srinivas et al. (2020) identified 13 threats and challenges impacting the river Ganga that included broad concerns ranging from biodiversity loss to open defecation near the river. To reduce anthropogenic induced threats and increase the overall health of the river basin, river rejuvenation and development must simultaneously address multiple social, economic, and environmental dimensions (Pal, 2023).
@@
In order to prioritize areas of the Ganga basin for actions such as conservation, restoration, intervention, and rejuvenation, it is crucial to perform vulnerability assessments. There are many vulnerability assessment methods, and a multitude of factors, dimensions, and subdimensions may be included in an assessment (Plummer et al., 2012). Critically, our assessment considers not only the immediate vulnerability of a given area but also the threat to the broader river system posed by potential sources of pollution in that area. This approach properly recognizes that the interaction of pollutants with local geophysical factors creates variation in environmental impacts downstream.
@@
The Analytical Hierarchical Process (AHP) is a multicriteria decision making process that combines pairwise comparisons and expert judgements to arrive at its conclusions. Developed by Saaty (1980), AHP has been used in numerous applications, ranging from resource allocation and site prioritization to risk assessment and conflict resolution. Climate and geophysical models have proven well-suited for AHP, due to the complexity of decision making under uncertainty with respect to factors impacting their features of interest. For example, Jhariya et al. (2021) used AHP in combination with Geospatial Information Systems (GIS), remote sensing, and vertical electrical soundings (VES) to identify potential zones of groundwater using weighted data on geology, geomorphology, rainfall, lineament, LULC (land use and land cover), drainage density, slope, soil type, and soil texture. Using AHP, they were able to estimate the likelihood of groundwater in each zone, with five levels of potential (low, medium, medium-high, high, and very high) with an accuracy of 80%.
1.1 AHP factors of waterway vulnerability and their Saaty rankings
@@
In this study, AHP was used as a characterization scheme rather than a decision making process; hence, factors and their relative rank order of importance were structured to highlight areas that may be vulnerable to pollution with an emphasis on urban pollution. Land use was identified using LULC, a widely recognized direct characterization of land types and includes development such as urbanization and agricultural uses. Urbanization may be directly related to major sources of contamination in water bodies, while agriculture is a major source of nonpoint source pollution including fertilizers and pesticides. Moreover, densely populated areas have a direct impact on surface water quality. For instance, population growth rate and water quality parameters, such as biochemical oxygen demand (BOD) and dissolved oxygen, were highly correlated in a Kelani river watershed in Sri Lanka (Liu et al., 2023). Given our emphasis on urban pollution and their direct impacts on vulnerability to pollution, population density (PD) was ranked as the most important factor (1 on the Saaty-scale), and LULC was ranked second with a Saaty-scale value of 2 in relative importance.
@@
Rainfall, slope, and drainage density indirectly impact vulnerability in that they are not sources of pollution, yet they directly influence erodibility and pollution mobility. For example, the runoff from steeper slopes was more likely to carry pollutants into streams than runoff from land use in flatter slopes (Yu et al., 2016), and effects on water quality originating from land use (agriculture, industrial, and residential) adjacent to water bodies were dependent on rainfall variability (Nobre et al., 2020). Drainage density (DD) is an indicator of surface runoff processes and direct runoff and pollution transport increase with greater DD (Bedient et al., 1978). Rainfall, DD, and slope were assigned Saaty-scale values of 4, 5, and 7 respectively because rainfall initiates pollutant movement events, slope is an important factor in pollutant transport, and DD influences final distribution of pollutants. Temperature is an environmental determinant of microbial activity and productivity including nuisance algal blooms. For instance, water surface temperature behavior in zones with algal bloom occurrences presented greater significant values, up to 3°C, than those with clearer water (Ferral et al., 2021). Moreover, seasonal increases of anaerobic bacteria in aquatic surface sediment during summer compared to cool months indicated a depletion of oxygen in the overlying water (Smoot and Findlay, 2001). Given the absence of high resolution site monitoring throughout the Ganga basin, land surface temperature (LST) was used as a surrogate temperature estimate. LST was assigned the lowest importance among all input factors (9 on the Saaty-scale) because it was expected to have a minor overall relative impact on vulnerability.
1.2 Potential Shortcomings of AHP
@@
While AHP offers numerous advantages in terms of decision making in remote sensing, Munier and Hontoria (2021) identified 30 potential shortcomings of the method. Given the importance of accurately understanding vulnerability, we reviewed all the shortcomings of AHP and assessed whether they were applicable for remote sensing data and our particular use case (Supplemental Table S1). For instance, the first critique “The Pair-Wise Method and Its Application in AHP” suggested that some problems may not be suited for AHP if their decision space is amorphous with many interrelated factors that have minimal differences in their relative importance. This critique was not applicable for our use-case, since our problem fits well to the AHP architecture. Similarly, the critique around rank reversals, which arise when uncertainty in the decision space is larger than the consistency of rank, was applicable on our use-case since a change in parameter weights might lead to an inaccurate assessment of the vulnerability of the river. To test for this and design a better method, we applied AHP while taking into account the effect of unknown variables (1-N AHP) and Fuzzy AHP as diagnostic analyses. To summarize, all the applicable shortcomings were addressed with the following approaches: Analytic Network Process (ANP) (Saaty, 2006), Fuzzy AHP (Saaty, 1982) , AHP with non-linear parametric influence (Nested AHP), 1-N AHP, and Fuzzy AHP while taking into account the effect of unknown factors (Fuzzy 1-N AHP) (Supplemental Table S1).
@@
In this paper, we apply AHP to our remote sensing datasets to assess the vulnerability of the river Ganga to pollution loads along a 1,330 km stretch of the river. We cover the problem formulation, results and inferences of AHP. We also include a deep-dive on each shortcoming of AHP in the context of Climate Modelling, addressing the list set forth by Munier and Hontoria (2021). We share whether each shortcoming was applicable to our problem statement, while suggesting and attempting alternative solutions to counteract the shortcomings. Through this dialogical process, we have arrived at two novel decision making approaches inspired by AHP, namely Nested AHP and 1-N AHP. The methodological contributions of this research are, at one level, these new variants of AHP we propose; and at another level, the demonstration of how incremental evaluation of the AHP process and its shortcomings relative to a given problem can reveal application-specific permutations of the original. Finally, we report the pollution vulnerability evaluation of the river Ganga produced by these processes.
2 Methods and Data
2.1 Study area and data source
@@
The area of analysis (26,609.4 km2) is a 20 km zone extending 10 km on both sides of the main channel of river Ganga from Haridwar, Uttarakhand, where it makes its debouche from the Siwalik Hills into the Indo-Gangetic plains to its confluence with river Ghaghara at Sitab Diara, Uttar Pradesh, India (Fig 1).
@@
The change in elevation along this 1,330 km stretch of Ganga is from 314 m to 76 m above mean sea level. Data from Landsat-8 (Level 2, Collection 2, Tier 1; pixel resolution, 30x30m) and other sources (Supplemental Table S2) were used to assess vulnerability in the area of analysis include slope, drainage density (DD), land surface temperature (LST), rainfall, land use/land cover (LULC), and population density (Fig 2). The drainage density and center line for the river Ganga were both derived from a digital elevation model (DEM) of the watershed using typical GIS hydrography processes, as previously described by Rahaman et al. (2015). Streams were extracted from NASADEM data as a vector format Shapefile using the R programming language, Whitebox, Terra, and sf packages. Lower order streams were removed to obtain the main Ganga river centerline with ArcGIS Pro. The Ganga center line down-river of the Uttar Pradesh border was removed, and the 20 km zone was established as a buffer around this shortened centerline within the river Ganga.
@@
Fig 1. This map shows the Ganga Basin, in gray, and the buffer region, in blue, which is the scope of our paper. The 20 km buffer zone extends 10 km on both sides of the main channel of river Ganga.
Fig 2. Visual representation of the data used in the calculation of AHP. AHP factors Population Density (a), Annual Rainfall (c), Drainage Density (d), Slope (e), and Land Surface Temperature (f) were divided into extreme low to extreme high using natural breaks. AHP factor LULC (b) was separated by land use class.
2.2 Analytical Hierarchy Process (AHP)
@@
To categorize the vulnerability of the area immediately around the river Ganga and to identify potential locations to implement our remediation strategies, we applied the AHP on Remote Sensing and other geospatial data as previously described by Jhariya et al. (2021). Environmental, topographical, and anthropogenic factors were subjectively ranked from equal importance to relatively greatest hierarchical importance across the full breadth of the 9-point Saaty scale based on their likelihood of impacting vulnerability (Section 1.1). A pairwise matrix constructed from these ranked factors and their reciprocal values were used to estimate a mean normalized weight for each factor (equation (1)). These normalized weights were applied to each of five factor classes that were identified by natural breaks in the data using the Jenks method (Table 1). LULC categories were numerically scored as vulnerability criteria based on their linkage to pollution. Lastly, overall vulnerability per pixel was determined from the sum of the factor vulnerability scores. Additional details describing factor rank orders, pairwise comparison matrices and their weights, as well as Jenks natural breaks are provided in Supplemental Information.
@@
Table 1: Factor classes used in Analytical Hierarchy Process decision matrix ranked by vulnerabilitya
@@
5: Extremely High >4036 Built-up >1410 >4.0 >35 >37
@@
1: Extremely Low <865 Water/Wetland <868 <0.6 <4 <25
aFactor classes were based on natural breaks (Jenks and Caspall, 1971).
@@ 2.2.1 Ranking method
$A=\begin{bmatrix}1&2&4&5&7&9\\ 0.5&1&3&4&5&7\\ 0.25&0.33&1&2&4&5\\ 0.2&0.25&0.5&1&3&5\\ 0.14&0.2&0.25&0.33&1&4\\ 0.11&0.14&0.2&0.2&0.25&1\end{bmatrix}$ (1)
@@
$\tilde{a}_{ij}=\frac{a_{ij}}{\sum_{i=1}^{n}a_{ij}}$ (2)
The priority vector $W$ (the eigenvector) was then derived by averaging the rows of the normalized matrix:
$w_{i}=\frac{1}{n}\sum_{j=1}^{n}\tilde{a}_{ij}$ (3)
@@
We then checked the consistency of the pairwise comparison matrix by calculating the Consistency Index (CI) and the Consistency Ratio (CR) (Saaty, 1980). The CI was calculated using the formula:
$CI=\frac{\lambda_{max}-n}{n-1}$ (4)
@@
$CR=CI/RI$ (5)
In our case, with a consistency ratio (CR) of 0.06, the matrix was considered consistent (as CR < 0.1). This consistency indicated that our pairwise comparison matrix was reliable for determining the weights. To create the vulnerability map of the Ganga buffer region, if
@@
    $W=\begin{bmatrix}w_{1}&w_{2}&w_{3}&w_{4}&w_{5}&w_{6}\end{bmatrix}$ (6)
@@
    $X(p)=\begin{bmatrix}x_{1}(p)&x_{2}(p)&x_{3}(p)&x_{4}(p)&x_{5}(p)&x_{6}(p)\end{bmatrix}$ (7)
We calculated the predicted vulnerability for each pixel by multiplying each factor’s parameter value by its corresponding AHP weight i.e. the dot product of the vector W with the transpose of the vector X(p):
$V(p)=W\cdot X(p)^{T}$ (8)
@@
2.3 Diagnostic Analyses
2.3.1 Nested AHP